# Lagrangian multiforms and dispersionless integrable systems

### v1

@@ 3.2 First heavenly equation
The Plebański first heavenly equation,
@@
Differentiation of $ab-pq=1$ implies $ab_{i}+ba_{i}=pq_{i}+qp_{i}$ which, by virtue of (25), gives the dispersion relation
@@
It remains to verify the identity $c^{i}_{j}=0$, i.e. that the $c_{i}$ are functions of $R^{i}$ only, which makes the above equations for $u$ and $\eta^{i}$ identical to equations (21). Thus, both first and second heavenly equations correspond to one and the same Lagrangian multiform (which is not surprising as both equations describe one and the same class of self-dual Ricci-flat metrics, and are Bäcklund-related).
@@ 4 Concluding remarks
  • We considered (integrable) PDEs $F=0$ with a (higher) conservation law $C$ with non-constant characteristic (cosymmetry) $S$, so that $\mathrm{d}C=SF$. Our examples suggest that the constrained system, $F=S=0$, is often involutive. It would be interesting to clarify what additional conditions are required for this to be the case.
@@
  • Our examples suggest that Gibbons-Tsarev type systems governing hydrodynamic reductions of linearly degenerate dispersionless integrable PDEs possess a Lagrangian multiform representation. Such systems are only of interest in dimensions $d\geq 4$ as, according to [20], Gibbons-Tsarev systems of three-dimensional linearly degenerate PDEs are linearisable. In contrast, Gibbons-Tsarev systems of higher-dimensional linearly degenerate PDEs are not linearisable (not even Darboux integrable).
@@ Acknowledgements
We thank S. Agafonov, V. Caudrelier, I. Krasilshchik, and M. Pavlov for useful discussions. MV is supported by the Engineering and Physical Sciences Research Council [Project reference EP/Y006712/1].
@@ A.1 Lagrangian multiform (4) and Euler-Lagrange equations
$B={u_{ijk}}-\frac{{u_{ij}}{u_{ik}}}{2{u_{i}}}-\frac{{u_{ij}}{u_{jk}}}{2{u_{j}}}-\frac{{u_{ik}}{u_{jk}}}{2{u_{k}}}.$
@@
If $A=0$, then $B=0$ is equivalent to the edge equation (5b),
$\frac{1}{2}a^{k}{\left(\frac{u_{ij}^{2}}{{u_{i}}u_{j}^{2}}-\frac{2{u_{iij}}}{{u_{i}}{u_{j}}}+\frac{2{u_{ii}}{u_{ij}}}{u_{i}^{2}{u_{j}}}\right)}+\frac{1}{2}a^{j}{\left(\frac{{u_{ik}}^{2}}{{u_{i}}{u_{k}}^{2}}-\frac{2{u_{iik}}}{{u_{i}}{u_{k}}}+\frac{2{u_{ii}}{u_{ik}}}{u_{i}^{2}{u_{k}}}\right)}=0,$
and equations obtained from this by cyclic permutation of $(i,j,k)$.
@@
$\frac{1}{u_{i}u_{j}u_{k}}\left(A_{i}-a^{i}u_{i}B-\frac{1}{2}\frac{u_{ij}}{u_{j}}A-\frac{1}{2}\frac{u_{ik}}{u_{k}}A\right).\qed$
@@
$A=0$ and $B=0$ together imply the planar equations (5a),
@@
Finally, we divide by $-u_{i}u_{j}u_{k}$ to find the planar Euler-Lagrange equation (5a). ∎

### v2

@@ 3.2 First heavenly equation
Plebański’s first heavenly equation,
@@
Differentiation of abpq=1ab-pq=1 implies $ab_{i}+ba_{i}=pq_{i}+qp_{i}$ which, by virtue of (25), gives the dispersion relation
@@
It remains to verify the identity $c^{i}_{j}=0$, i.e. that the $c_{i}$ are functions of $R^{i}$ only, which makes the above equations for $u$ and $\eta^{i}$ identical to equations (21). Thus, Gibbons-Tsarev systems governing hydrodynamic reductions of the first and second heavenly equations come from one and the same Lagrangian multiform (which is not surprising as both equations describe one and the same class of self-dual Ricci-flat metrics and are Bäcklund-related [31]).
@@ 4 Concluding remarks
  • We considered (integrable) PDEs $F=0$ with a (higher) conservation law $C$ with non-constant characteristic (cosymmetry) $S$, so that $\mathrm{d}C=FS$. Our examples suggest that the constrained system, $F=S=0$, is often involutive. It would be interesting to clarify what additional conditions are required for this to be the case.
@@
  • Our examples suggest that Gibbons-Tsarev systems, governing hydrodynamic reductions of linearly degenerate dispersionless integrable PDEs, possess a Lagrangian multiform representation.
@@ Acknowledgements
We thank S. Agafonov, V. Caudrelier, I. Krasilshchik, O. Morozov, V. Novikov, M. Pavlov, and Y. Suris for clarifying discussions. We also thank the reviewers for useful comments.
MV is supported by the Engineering and Physical Sciences Research Council [Project reference EP/Y006712/1].
Declarations
Conflict of interest
There is no conflict of interest related to this article.
Data availability
A pre-print article has appeared as arXiv:2503.22615. There is no additional data related to this article.
@@ A.1 Lagrangian multiform (4) and Euler-Lagrange equations
$B=2u_{ijk}-\frac{{u_{ij}}{u_{ik}}}{{u_{i}}}-\frac{{u_{ij}}{u_{jk}}}{{u_{j}}}-\frac{{u_{ik}}{u_{jk}}}{{u_{k}}}.$
@@
If $A=0$ and its differential consequences hold, then $B=0$ is equivalent to the edge equation (5b).
@@
$\frac{1}{u_{i}u_{j}u_{k}}\left(2A_{i}-a^{i}u_{i}B-\frac{u_{ij}}{u_{j}}A-\frac{u_{ik}}{u_{k}}A-2\frac{u_{ii}}{u_{i}}A\right).\qed$
@@
$A=0$ and $B=0$, together with their differential consequences, imply the planar equations (5a),
@@
Finally, we divide by $u_{i}u_{j}u_{k}$ to find the planar Euler-Lagrange equation (5a). ∎
@@
Analogous calculations show that relations (9) and (7) are equivalent to the multiform Euler-Lagrange equations in the translationally non-invariant case (8). The multiform Euler-Lagrange equations in this case are
$(x^{i}-x^{j})\left(\left(\frac{u_{ij}^{2}}{u_{i}^{2}u_{j}}\right)_{i}+\left(\frac{u_{ij}^{2}}{u_{i}u_{j}^{2}}\right)_{j}+\left(\frac{2u_{ij}}{u_{i}u_{j}}\right)_{ij}\right)$
$\hskip 85.35826pt+\frac{u_{ij}^{2}}{u_{i}^{2}u_{j}}-\frac{u_{ij}^{2}}{u_{i}u_{j}^{2}}+\left(\frac{2u_{ij}}{u_{i}u_{j}}\right)_{j}-\left(\frac{2u_{ij}}{u_{i}u_{j}}\right)_{i}=0,$ (38a)
$(x^{j}-x^{i})\left(\frac{u_{ij}^{2}}{u_{i}u_{j}^{2}}+2\frac{u_{ii}u_{ij}}{u_{i}^{2}u_{j}}-2\frac{u_{iij}}{u_{i}u_{j}}\right)$
$\hskip 85.35826pt-(x^{k}-x^{i})\left(\frac{u_{ik}^{2}}{u_{i}u_{k}^{2}}+2\frac{u_{ii}u_{ik}}{u_{i}^{2}u_{k}}-2\frac{u_{iik}}{u_{i}u_{k}}\right)+\frac{2u_{ij}}{u_{i}u_{j}}-\frac{2u_{ik}}{u_{i}u_{k}}=0,$ (38b)
$(x^{i}-x^{j})\frac{u_{ij}}{u_{i}u_{j}}+(x^{j}-x^{k})\frac{u_{jk}}{u_{j}u_{k}}+(x^{k}-x^{i})\frac{u_{ik}}{u_{i}u_{k}}=0.$ (38c)
The proof of Lemma 2 does not change and the proof of Lemma 3 requires one additional step of adding $\frac{B}{2u_{j}u_{k}}-\frac{B}{2u_{i}u_{j}}$.